# Nonharmonic multivariate Fourier transforms and matrices: condition numbers and hyperplane geometry

### v1

@@ 6 Proofs of propositions and lemmas
Fix any nonzero $u\in\mathbb{C}^{s}$. We start with $\mathcal{F}_{\Omega,\mathcal{X}}$. We have
@@
Now we move onto $\Phi_{\Omega,\mathcal{X}}$. A similar calculation, together with the Poisson summation formula and the assumed decay condition on $\widehat{\psi}$, yields
$|\Phi_{\Omega,\mathcal{X}}u|_{2}^{2}$ $=\sum_{\omega\in\Omega\cap\mathbb{Z}^{d}}\Big{|}\sum_{k=1}^{s}u_{k}e^{-2\pi i\omega\cdot x_{k}}\Big{|}^{2}\geq\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)\Big{|}\sum_{k=1}^{s}u_{k}e^{-2\pi i\omega\cdot x_{k}}\Big{|}^{2}$
$=\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)e^{2\pi i\omega\cdot(x_{j}-x_{k})}=\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\sum_{\omega\in\mathbb{Z}^{d}}\psi(\omega+x_{j}-x_{k}).$
This switch of summation is justified by Tonelli’s theorem, since $\widehat{\psi}(\omega)\leq 0$ for all $\omega\not\in\Omega$ and $\Omega$ is assumed to be compact. Notice that if $j\not=k$, then for all $\omega\in\mathbb{Z}^{d}$, we have $|\omega+x_{j}-x_{k}|_{p}\geq|x_{j}-x_{k}|_{\ell^{p}(\mathbb{T}^{d})}\geq\Delta_{p}(\mathcal{X})>\delta.$ This implies $\psi(\omega+x_{j}-x_{k})=0$ for all $\omega\in\mathbb{Z}^{d}$ whenever $j\not=k$ because $\psi$ is supported in $\Omega_{\delta}^{p}$. Thus, for any unit norm $u\in\mathbb{C}^{s}$, we have
@@
This shows that $\sigma_{\min}(\Phi_{\Omega,\mathcal{X}})\geq\sqrt{\psi(0)}$. For the upper bound on $\sigma_{\max}(\Phi_{\Omega,\mathcal{X}})$, the argument is analogous, except we use $\varphi$ instead and make appropriate adjustments to the above inequalities.
@@
On the other hand, using that $f\in\mathcal{B}(\Omega)$ and Cauchy-Schwarz, we have
@@
$|1-e^{2\pi iq\cdot u}|=2\pi|q\cdot u|\,{\rm sinc}(|q\cdot u|)\geq\frac{\pi}{2\alpha}|u|_{p^{\prime}}\,{\rm sinc}\Big{(}\frac{|u|_{p^{\prime}}}{4\alpha}\Big{)}\geq\frac{\pi}{2\alpha}|u|_{p^{\prime}}\,{\rm sinc}\Big{(}\frac{1}{4}\Big{)}=\frac{\sqrt{2}}{\alpha}|u|_{p^{\prime}},$
@@
Let $\theta_{k}\in\mathbb{R}^{d}$ be a unit $\ell^{2}$ norm vector orthogonal to $\mathcal{H}_{k}$. Since $\eta_{k}$ is the $\ell^{2}$ distance between $\mathcal{H}_{k}$ and 00, we set $u_{k}:=\eta_{k}\theta_{k}\in\mathcal{H}_{k}$. Next, we consider the function $h\in L^{\infty}(\mathbb{R}^{d})$ by
@@
where the last inequality follows by assumption. Consequently, $h$ is well defined. Second, is constant on hyperplanes orthogonal to $\mathcal{H}_{k}$, and in particular, $h(0)=1$ and $h$ vanishes on $\mathcal{H}_{k}$. Finally, in the sense of distributions, the Fourier transform of $h$ is supported in the set
$\Big{\{}\frac{n}{r+1}\theta_{1},\frac{n}{r+1}\theta_{2},\,\dots,\,\frac{n}{r+1}\theta_{r}\Big{\}}.$ (6.8)
@@
$|1-e^{2\pi iu_{k}\cdot n\theta_{k}/(r+1)}|=|1-e^{2\pi in\eta_{k}/(r+1)}|\geq\frac{2\pi n\eta_{k}}{r+1}\,{\rm sinc}\Big{(}\frac{1}{4}\Big{)}=\frac{4\sqrt{2}n\eta_{k}}{r+1}.$
@@
Let $\theta_{k}\in\mathbb{R}^{d}$ be a unit $\ell^{2}$ norm vector orthogonal to $\mathcal{H}_{k}$. Since $\eta_{k}$ is the $\ell^{2}$ distance between $\mathcal{H}_{k}$ and 00, we set $u_{k}:=\eta_{k}\theta_{k}\in\mathcal{H}_{k}$. Consider the trigonometric polynomial $h$, where
@@
The denominators of this function are all nonzero because $\theta_{k}$ is parallel to $q_{k}$, which establishes that $q_{k}\cdot u_{k}=\eta_{k}|q_{k}|_{2}>0$. To control the denominators, notice that $\eta_{k}|q_{k}|_{2}\leq\eta_{k}\frac{n}{r+1}\leq\frac{1}{4}$, so
$|1-e^{2\pi iq_{k}\cdot u_{k}}|\geq 2\pi\eta_{k}|q_{k}|_{2}{\rm sinc}\Big{(}\frac{1}{4}\Big{)}=4\sqrt{2}\eta_{k}|q_{k}|_{2}.$
@@ Conclusion
One could consider sampling the Fourier transform from $TB_{m}$ or $TQ_{m}$ for some transformation $T$ on $\mathbb{R}^{d}$. This would correspond to using a (possibly rotated) ellipsoidal or parallelpiped aperture, respectively. In this case, our results for the continuous Fourier operator readily generalize due to simple transform properties – one would just replace the $\ell^{p}$ distance on $\mathcal{X}$ with the $\ell^{p}$ metric in the transformed space, $|x|_{p,T}:=|Tx|_{p}$. For Fourier matrices, we expect exactly the same effect. However, it does not have nice transformation properties like its continuous counterpart (e.g., $B_{m}$ is closed under rotations while $B_{m}\cap\mathbb{Z}^{d}$ is not), so we cannot simply perform a change of variables – one would need to modify some of the lemmas and arguments in this paper.

### v2

@@ 6 Proofs of propositions and lemmas
Fix any nonzero $u\in\mathbb{C}^{s}$. We start with $\mathcal{F}_{\Omega,\mathcal{X}}$. Since $\widehat{\psi}$ is integrable, we have
@@
Now we move onto $\Phi_{\Omega,\mathcal{X}}$. A formal calculation with the Poisson summation formula yields
$\begin{split}&|\Phi_{\Omega,\mathcal{X}}u|_{2}^{2}=\sum_{\omega\in\Omega\cap\mathbb{Z}^{d}}\Big{|}\sum_{k=1}^{s}u_{k}e^{-2\pi i\omega\cdot x_{k}}\Big{|}^{2}\geq\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)\Big{|}\sum_{k=1}^{s}u_{k}e^{-2\pi i\omega\cdot x_{k}}\Big{|}^{2}\\ &=\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)e^{2\pi i\omega\cdot(x_{j}-x_{k})}=\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\sum_{\omega\in\mathbb{Z}^{d}}\psi(\omega+x_{j}-x_{k})=\psi(0)|u|_{2}^{2}.\end{split}$ (6.1)
We will provide a rigorous derivation at the end, which requires justifying the switch of summation and Poisson summation formula. Thus, for any unit norm $u\in\mathbb{C}^{s}$, we have
@@
This shows that $\sigma_{\min}(\Phi_{\Omega,\mathcal{X}})\geq\sqrt{\psi(0)}$. For the upper bound on $\sigma_{\max}(\Phi_{\Omega,\mathcal{X}})$, the argument is analogous, except we use $\varphi$ instead and make appropriate adjustments.
Let us return back to the justification of (6.1), which are based on standard ideas such as in [30, Chapter VI, Section 1.15]. Let $\eta\colon\mathbb{R}^{d}\to\mathbb{R}$ be an infinitely differentiable function that is compactly supported in $Q_{1/4}$, $0\leq\widehat{\eta}\leq 1$, and $\widehat{\eta}=1$ on $\Omega$. For any $\varepsilon\in(0,1)$, consider the function $\eta_{\varepsilon}(x):=\varepsilon^{-d}\eta(x/\varepsilon)$. Then $\widehat{\eta}_{\varepsilon}=\eta(\varepsilon\omega)$ and in particular, $\widehat{\eta}_{\varepsilon}=1$ on $\Omega$ for all $\varepsilon\in(0,1)$. Note $\widehat{\psi}$ is bounded since $\psi$ is continuous and supported in a compact set, and $\widehat{\eta}$ decays faster than any polynomial due to $\eta$ being infinitely differentiable. Hence $\{\widehat{\psi}(\omega)\widehat{\eta}(\varepsilon\omega)\}_{\omega\in\mathbb{Z}^{d}}\in\ell^{1}(\mathbb{Z}^{d})$, which enables us to switch the order of summation in the following calculation,
$\begin{split}|\Phi_{\Omega,\mathcal{X}}u|_{2}^{2}&=\sum_{\omega\in\Omega\cap\mathbb{Z}^{d}}\Big{|}\sum_{k=1}^{s}u_{k}e^{-2\pi i\omega\cdot x_{k}}\Big{|}^{2}\\ &\geq\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)\widehat{\eta}(\varepsilon\omega)\Big{|}\sum_{k=1}^{s}u_{k}e^{-2\pi i\omega\cdot x_{k}}\Big{|}^{2}\\ &=\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)\widehat{\eta}(\varepsilon\omega)e^{2\pi i\omega\cdot(x_{j}-x_{k})}.\end{split}$ (6.2)
Due the assumption that $\psi$ is supported in $\Omega_{\delta}^{p}$ for some $\delta\in(0,\frac{1}{2})$, we see $\psi$ can be interpreted as a continuous function on $\mathbb{T}^{d}$ and that $\widehat{\psi}(\omega)$ is the Fourier coefficient of $\psi$ at $\omega\in\mathbb{Z}^{d}$. Likewise, $\eta_{\varepsilon}$ is compactly supported in $Q_{\varepsilon/4}$, so it is also an infinitely differentiable function on $\mathbb{T}^{d}$ and $\widehat{\eta}(\varepsilon\omega)$ is the Fourier coefficient of $\eta_{\varepsilon}$ at $\omega\in\mathbb{Z}^{d}$. Thus, the sum over $\omega\in\mathbb{Z}^{d}$ on the right hand side of (6.2) is the Fourier series of $\psi*\eta_{\varepsilon}$ (treated as a periodic function) evaluated at $x_{j}-x_{k}$. This Fourier series converges pointwise due to $\{\widehat{\psi}(\omega)\widehat{\eta}(\varepsilon\omega)\}_{\omega\in\mathbb{Z}^{d}}\in\ell^{1}(\mathbb{Z}^{d})$, and so
$\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\sum_{\omega\in\mathbb{Z}^{d}}\widehat{\psi}(\omega)\widehat{\eta}(\varepsilon\omega)e^{2\pi i\omega\cdot(x_{j}-x_{k})}=\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}(\psi*\eta_{\varepsilon})(x_{j}-x_{k}).$
Combining the above observations yields the inequality
$|\Phi_{\Omega,\mathcal{X}}u|_{2}^{2}\geq\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}(\psi*\eta_{\varepsilon})(x_{j}-x_{k}).$
Now we are in position to take the limit as $\varepsilon\to 0$. By assumption, $\eta$ is integrable and $\int_{\mathbb{T}^{d}}\eta_{\varepsilon}=\int_{\mathbb{T}^{d}}\eta=\widehat{\eta}(0)=1$. Then $\{\eta_{\varepsilon}\}_{\varepsilon>0}$ is an approximation to the identity as $\varepsilon\to 0$. Hence, $\psi*\eta_{\varepsilon}\to\psi$ pointwise since $\psi$ is assumed to be continuous. Therefore, taking the limit now shows that
$|\Phi_{\Omega,\mathcal{X}}u|_{2}^{2}\geq\sum_{j=1}^{s}\sum_{k=1}^{s}\overline{u_{j}}u_{k}\psi(x_{j}-x_{k})=\sum_{j=1}^{s}|u_{j}|^{2}\psi(0)=\psi(0)|u|^{2}_{2}.$
This now proves inequality (6.1).
@@
On the other hand, using that $f\in\mathcal{W}(\Omega)$ and Cauchy-Schwarz, we have
@@
$|1-e^{2\pi iq\cdot u}|=2\pi|q\cdot u|\,{\rm sinc}(|q\cdot u|)\geq\frac{\pi}{2\alpha}|u|_{p^{\prime}}\,{\rm sinc}\left(\frac{|u|_{p^{\prime}}}{4\alpha}\right)\geq\frac{\pi}{2\alpha}|u|_{p^{\prime}}\,{\rm sinc}\left(\frac{1}{4}\right)=\frac{\sqrt{2}}{\alpha}|u|_{p^{\prime}},$
@@
Enumerate the $r$ hyperplanes by $\mathcal{H}_{1},\dots,\mathcal{H}_{r}$. Since $\eta_{k}$ is the $\ell^{2}$ distance between $\mathcal{H}_{k}$ and 00, we can find a unit $\ell^{2}$ norm vector $\theta_{k}\in\mathbb{R}^{d}$ that is orthogonal to $\mathcal{H}_{k}$ and $u_{k}:=\eta_{k}\theta_{k}\in\mathcal{H}_{k}$. Next, we consider the function $h\in L^{\infty}(\mathbb{R}^{d})$ by
@@
where the last inequality follows by assumption. Consequently, $h$ is well defined. Second, the $k$-th factor in the definition of $h$ is constant on hyperplanes parallel to $\mathcal{H}_{k}$, i.e., orthogonal to $\theta_{k}$. Indeed, for any $x\in\mathbb{R}^{d}$ and $v\in\mathbb{R}^{d}$ that is orthogonal to $\theta_{k}$, we see that $(x+v)\cdot\theta_{k}=x\cdot\theta_{k}$. Since the $k$-th factor is zero when evaluated at $u_{k}\in\mathcal{H}_{k}$, we see that the $k$-th factor is zero on $\mathcal{H}_{k}$. Thus, $h(0)=1$ and $h$ vanishes on $\mathcal{H}_{1}\cup\cdots\cup\mathcal{H}_{r}$. Finally, in the sense of distributions, the Fourier transform of $h$ is supported in the set
$\left\{\frac{n}{r+1}\theta_{1},\,\frac{n}{r+1}\theta_{2},\,\,\dots,\,\frac{n}{r+1}\theta_{r}\right\}.$ (6.10)
@@
$|1-e^{2\pi iu_{k}\cdot n\theta_{k}/(r+1)}|=|1-e^{2\pi in\eta_{k}/(r+1)}|\geq\frac{2\pi n\eta_{k}}{r+1}\,{\rm sinc}\left(\frac{1}{4}\right)=\frac{4\sqrt{2}n\eta_{k}}{r+1}.$
@@
Enumerate the $r$ hyperplanes by $\mathcal{H}_{1},\dots,\mathcal{H}_{r}$. Since $\eta_{k}$ is the $\ell^{2}$ distance between $\mathcal{H}_{k}$ and 00, we can find a unit $\ell^{2}$ norm vector $\theta_{k}\in\mathbb{R}^{d}$ that is orthogonal to $\mathcal{H}_{k}$ and $u_{k}:=\eta_{k}\theta_{k}\in\mathcal{H}_{k}$. Consider the trigonometric polynomial $h$, where
@@
The denominators of this function are all nonzero because $\theta_{k}$ is parallel to $q_{k}$, which establishes that $|q_{k}\cdot u_{k}|=|q_{k}|_{2}|\theta_{k}\cdot u_{k}|=\eta_{k}|q_{k}|_{2}>0$. To control the denominators, notice that $\eta_{k}|q_{k}|_{2}\leq\eta_{k}\frac{n}{r+1}\leq\frac{1}{4}$, so
$|1-e^{2\pi iq_{k}\cdot u_{k}}|\geq 2\pi\eta_{k}|q_{k}|_{2}{\rm sinc}\left(\frac{1}{4}\right)=4\sqrt{2}\eta_{k}|q_{k}|_{2}.$
@@ Conclusion
One could consider sampling the Fourier transform from $TB_{m}$ or $TQ_{m}$ for some invertible linear transformation $T$ on $\mathbb{R}^{d}$. This would correspond to using a (possibly rotated) ellipsoidal or parallelpiped aperture, respectively. In this case, our results for the continuous Fourier operator readily generalize due to simple transform properties – one would just replace the $\ell^{p}$ distance on $\mathcal{X}$ with the $\ell^{p}$ metric in the transformed space, $|x|_{p,T}:=|Tx|_{p}$. For Fourier matrices, we expect exactly the same effect. However, it does not have nice transformation properties like its continuous counterpart (e.g., $B_{m}$ is closed under rotations while $B_{m}\cap\mathbb{Z}^{d}$ is not), so we cannot simply perform a change of variables – one would need to modify some of the lemmas and arguments in this paper.